OpenStudy (anonymous):

Anita and Joelle bowl together and their combined total score for one game was 425 points. Anita’s score was 40 less than twice Joelle’s. What were their scores?

OpenStudy (anonymous):

set variables for the bowlers to make it easy, let's call Anita A and Joelle J. A=2J-40 425=3J-40 ... solve for J to get J = 155 now plug in 155 for J in A= 2J-40 to get A = 2(155)-40 = 270 So Anita scored 270 and Joelle scored 155... you can double check this by doubling Joelle's score of 155 and subtracting 40... does it equal 270? yes.

OpenStudy (anonymous):

Thanks!